# LIR: A Lightweight Baseline for Image Restoration

### v2

@@ 3.3 Attention Block
While different types of filters in Adaptive Filter (AF) enable LAA blocks to extract rich information, AF alone is not enough, as we need to take full advantage of the rich information extracted by AF. Thus, a strong and efficient Attention Block is introduced, which consists of four components: ResBlock, Patch Attention (PA), SimpleGate, and ResCABlock. We first use the ResBlock to preprocess input features, followed by the extraction of global information using our PA module. The efficacy of the gate function has been demonstrated in prior research [29, 52, 4]. Here, we simply employ the SimpleGate module from [4] to further extract dependencies between features. Finally, we leverage the ResCABlock to extract channel-wise information.
3.3.1 ResBlock and ResCABlock
@@
Both ResBlock and ResCABlock are up-down style residual blocks that include a learnable parameter in the trunk. The only difference between ResBlock and ResCABlock is that the latter has a channel attention module at the tail. The Channel Attention module [68], based on Squeeze-and-Excitation (SE) [17], has gained much popularity for its ability to extract channel-wise information. Many studies [27, 6, 4, 69] have directly utilized this module in their models. Following them, in our Attention Block, we also incorporate the Channel Attention module into the ResBlock to extract channel-wise information. In addition, the structure of ResBlock and ResCABlock are designed to be simple to meet lightweight and efficient requirements.
[…]
[width=]figures/patch_attention_2.eps
Figure 4: The flow of the proposed Patch Attention.
3.3.2 Patch Attention
@@
To give the Attention Block the ability to capture global information, we introduce the Patch Attention (PA) (Fig. 4). In the self-attention part of the transformer, the attention map is calculated using two matrix multiplications and a softmax operation between q, k, and v, which incurs significant computational overhead. Instead, we simplify the attention map computation in our PA by using just one MLP and one element-wise multiplication.
@@
Table 1: Image Deraining results evaluated by PSNR and SSIM on Rain100L dataset. Where $\triangle$ stands for the official code that is unavailable or not based on PyTorch.
[…]
0.9 Method Type Rain100L [56] Params [M] GFLOPs PSNR $\downarrow$ SSIM UMRL [57] CNN 29.18 0.923 0.98 9.31 PReNet [42] RNN 32.44 0.950 0.17 37.27 AirNet [24] CNN 34.90 0.966 8.93 169.47 MPRNet [59] CNN 36.40 0.965 20.13 960.39 SPAIR [40] CNN 36.93 0.969 $\triangle$ $\triangle$ HINet [5] CNN 37.28 0.970 88.67 95.92 MAXIM-2S [52] MLP 38.06 0.977 22.50 80.13 SFNet [7] CNN 38.21 0.974 13.27 70.00 DRCNet [25] CNN 38.23 0.976 18.90 $\triangle$ Restormer [58] Transformer 38.99 0.978 26.13 79.31 LIR (Ours) CNN 38.96 0.983 7.31 63.96
[…]
[width=]figures/derain100l.eps
Figure 5: The visual comparisons for LIR, Restormer [58], SFNet [7], and HINet [5]. LIR produces state-of-the-art visual results.
4 Experiment
4.1 Experiment Settings
We set the number of LAA Blocks in the LIR to [3,3,3,3,4], and the width to 48. During the training, we use the $L_{1}$ loss function, AdamW ($\beta_{1}$=0.9, $\beta_{2}$=0.999) optimizer with an initial learning rate of $3e^{-4}$, and set weight decay as $1e^{-4}$. Random image cropping and horizontal flipping are used as data augmentation strategies. Following [58], we use the cosine annealing strategy to gradually decrease the learning rate from $3e^{-4}$ to $1e^{-6}$, and use the progressive learning, where training patch size and batch size pairs are: [($128^{2}$,32), ($160^{2}$,16), ($256^{2}$,8)] at iterations [250K, 250K, 250K] for deblurring, deraining, and denoising. For dehazing, we set [($128^{2}$,32), ($160^{2}$,16)] at iterations [350K, 400K]. The number of parameters and computation for each method are calculated based on the input size of (3,192,192). Specifically, parameters are counted using the PyTorch, and the computation is measured using the thop package. All experiments are conducted using the official release code for each method. For methods that do not have released code or are not based on PyTorch, we have marked them with $\triangle$ in the tables.
[…]
[width=0.6]figures/dehaze.eps
@@ 4.1 Experiment Settings
Figure 6: The visual comparisons for LIR, FFA-Net [41], GridDehazeNet [30], and AirNet [24]. LIR produces better visual results.
4.2 Image Deraining
The Rain14000 [15] (11,200 training images) and Rain100H [56] (1,254 training images) are used to train LIR, and the Rain100L [56] datasets is used for evaluation. The results are presented in Tab. 1. In the Rain100L results (Tab. 1), LIR with minimal parameters and computation outperforms the DRCNet [25] and achieves comparable performance to Restormer [58] with only a slight 0.02 dB decrease on PSNR. In addition, our LIR achieves state-of-the-art on SSIM, which significantly outperforms Restormer. As GFLOPs do not accurately reflect the computations of the transformer model, the actual computation of LIR is much lower compared to Restormer, as shown in Tab. 1.
It is worth noting that with fewer parameters and computations, LIR significantly outperforms SFNet [7], the same lightweight design method as ours. Fig. 5 shows visual comparisons among LIR, Restormer, SFNet, and HINet. It can be demonstrated that LIR produces state-of-the-art visual results, despite LIR having a slightly lower PSNR than Restormer.
Table 2: Image dehazing results evaluated by PSNR and SSIM on SOTS outdoor dataset. Where $\triangle$ stands for the official code that is unavailable or not based on PyTorch.
[…]
0.9 Method Type SOTS outdoor [23] Params [M] GFLOPs PSNR $\downarrow$ SSIM FD-GAN [11] GAN 23.15 0.920 13.98 38.90 AirNet [24] CNN 23.18 0.900 8.93 169.47 IDR [62] CNN 25.24 0.943 15.34 $\triangle$ GridDehazeNet [30] CNN 30.86 0.982 0.96 12.05 U2-Former [20] Transformer 31.10 0.976 $\triangle$ $\triangle$ FFA-Net [41] CNN 33.38 0.980 4.46 161.74 MAXIM-2S [52] MLP 34.19 0.985 22.5 80.13 DehazeFormer [44] Transformer 34.95 0.984 2.52 13.37 LIR (Ours) CNN 34.85 0.985 7.31 63.96
[…]
[width=]figures/denoise.eps
Figure 7: The visual comparisons for LIR, Restormer [58], AirNet [24], and MPRNet [59]. LIR produces state-of-the-art visual results.
4.3 Image Dehazing
The OTS [23] (72,315 training images) is used to train LIR, and the Synthetic Objective Testing Set (SOTS) outdoor [23] is used for evaluation. In the SOTS outdoor results (Tab. 2), LIR significantly outperforms both the MAXIM-2S [52] (MLP-based) and the U2-Former [20] (transformer-based) with improvement in PSNR of 0.66 dB and 3.75 dB. In addition, LIR achieves comparable performance on PSNR and better performance on SSIM to DehazeFormer. It is worth noting that GFLOPs and parameters do not accurately reflect the complexity of the transformer model, the actual complexity of LIR is much lower compared to DehazeFormer. Visual comparisons are shown in the Fig. 6.
@@ 4.3 Image Dehazing
Table 3: Image denoising results evaluated by PSNR and SSIM on CBSD68 and Urban100 datasets. Where $\triangle$ stands for the official code that is unavailable or not based on PyTorch and light blue for the method that is learned to handle various noise levels.
[…]
0.9 Method Type CBSD68 [35] $\downarrow$ Urban100 [18] Params [M] GFLOPs $\sigma=15$ $\sigma=25$ $\sigma=50$ $\sigma=15$ $\sigma=25$ $\sigma=50$ FD-GAN [11] GAN 30.25 28.81 26.43 - - - 13.98 38.90 DL [14] CNN 33.05 30.41 26.90 - - - $\triangle$ $\triangle$ MPRNet [59] CNN 33.54 30.89 27.56 - - - 20.13 960.39 AirNet [24] CNN 33.92 31.26 28.00 34.40 32.10 28.88 8.93 169.47 BRDNet [46] CNN 34.10 31.43 28.16 34.42 31.99 28.56 $\triangle$ $\triangle$ IDR [62] CNN 34.11 31.60 28.14 33.82 31.29 28.07 15.34 $\triangle$ DRUNet [64] CNN 34.30 31.69 28.51 34.81 32.60 29.61 32.64 80.78 Restormer [4] Transformer 34.39 31.78 28.59 35.06 32.91 30.02 26.13 79.31 LIR (Ours) CNN 34.31 31.69 28.50 34.81 32.59 29.57 7.31 63.96
4.4 Image Denoising
We perform denoising experiments on synthetic benchmark datasets generated by additive white Gaussian noise. The DFWB (DIV2K [1], Flickr2K, WED [34], BSD500 [2]) datasets with 8,594 training images are used to train LIR. For evaluation, we use the CBSD68 [35] and Urban100 [18] datasets. The light blue color in the Tab. 3 indicates that the method is trained using a blind denoising setting (random $\sigma$ from [0,50]) and then evaluated under a fixed $\sigma$ (0, 25, or 50). Other methods are trained with a fixed $\sigma$ and evaluated with the same $\sigma$ as training. As shown in the Tab. 3, LIR achieves comparable performance to Restormer [58] and significantly outperforms all other models on the CBSD68 dataset with a blind denoising setting. Moreover, it is worth noting that with a 77% reduction in parameters and a 20% reduction in computations, our LIR achieves comparable performance to DRUNet [64] on Urban100, while the DRUNet is trained under fixed $\sigma$ setting and our LIR is under random $\sigma$ setting. The visual comparisons between LIR, Restormer, AirNet, and MPRNet are shown in Fig. 7. Although LIR is not as high as Restormer on PSNR, the state-of-the-art visual result is achieved.
@@ 4.4 Image Denoising
Table 4: Image deblurring results evaluated by PSNR and SSIM on GoPro and HIDE datasets. Where $\triangle$ stands for the official code that is unavailable or not based on PyTorch.
[…]
0.9 Method Type GoPro [36] HIDE [43] Params [M] GFLOPs PSNR $\downarrow$ SSIM PSNR $\downarrow$ SSIM IDR [62] CNN 27.87 0.846 - - 15.34 $\triangle$ DBGAN [65] GAN 31.10 0.942 28.94 0.915 11.58 189.71 MT-RNN [37] RNN 31.15 0.945 29.15 0.918 $\triangle$ $\triangle$ DMPHN [60] CNN 31.20 0.940 29.09 0.924 21.70 $\triangle$ Zhe et al. [21] CNN 31.79 0.949 - - $\triangle$ $\triangle$ Suin et al. [45] CNN 31.85 0.948 29.98 0.939 $\triangle$ $\triangle$ SPAIR [40] CNN 32.06 0.950 30.29 0.931 $\triangle$ $\triangle$ LIR (Ours) CNN 32.19 0.931 30.31 0.917 7.31 63.96
4.5 Image Deblurring
@@
The GoPro [36] (2,103 training images) dataset is used to train LIR, and the GoPro (1,111 images) and HIDE [43] (2,025 images) datasets are used for evaluation. Tab. 4 demonstrates that our LIR significantly outperforms other methods with fewer parameters and computations. Specifically, LIR is 0.13 dB higher than SPAIR [40], and 0.34 dB significantly higher than Suin et al [45].
4.6 Ablation Studies
@@
We conduct ablation experiments on the Rain100L dataset, using unchanged LIR as the baseline. The ablation experiments are performed on three types: Simple Residual, No Patch Attention, and No Adaptive Filter. Simple Residual refers to using vanilla global and local residual connections in the model, without removing the degradation in residual connections through structural design. In addition, No Adaptive Filter refers to using a vanilla convolution to replace the Adaptive Filter for fair comparison. Therefore, the number of parameters is unchanged. The results in the Tab. 5 demonstrate that the absence of the Adaptive Filter has the greatest negative impact on LIR. In addition, the experiment without Patch Attention demonstrates that our proposed Patch Attention module is efficient and computation-friendly, only bringing 0.8 M parameters and 1.81 GFLOPs computations.
Table 5: Ablation stuides are performed on Rain100L dataset.
Metric Baseline Simple Residual Patch Attention Adaptive Filter
PSNR 38.96 38.87 38.85 38.66
Params [M] 7.31 6.86 6.51 7.31
GFLOPs 63.96 49.15 62.15 62.40
5 Conclusion and Limitation
5.1 Conclusion

### v3

@@ 3.3. Attention Block
While different types of filters in Adaptive Filter enable LAA blocks to extract rich information, Adaptive Filter alone is not enough, as we need to take full advantage of the rich information extracted by Adaptive Filter. Thus, a strong and efficient Attention Block is introduced, which consists of four components: ResBlock, Patch Attention (PA), SimpleGate, and ResCABlock. We first use the ResBlock to preprocess input features, followed by the extraction of global information using our PA module. The efficacy of the gate function has been demonstrated in prior research (Liu et al., 2021a; Tu et al., 2022; Chen et al., 2022a). Here, we simply employ the SimpleGate module from (Chen et al., 2022a) to further extract dependencies between features. Finally, we leverage the ResCABlock to extract channel-wise information.
Table 1. Image Deraining results evaluated by PSNR and SSIM on Rain100L dataset.
Method Type Rain100L (Yang et al., 2017) Params [M] GFLOPs
PSNR SSIM
UMRL (Yasarla and Patel, 2019) CNN 29.18 0.923 0.98 9.31
PReNet (Ren et al., 2019) RNN 32.44 0.950 0.17 37.27
AirNet (Li et al., 2022a) CNN 34.90 0.966 8.93 169.47
MPRNet (Zamir et al., 2021) CNN 36.40 0.965 20.13 960.39
SPAIR (Purohit et al., 2021) CNN 36.93 0.969 - -
HINet (Chen et al., 2021a) CNN 37.28 0.970 88.67 95.92
IRNeXt (Cui et al., 2023) CNN 38.14 0.972 5.46 23.60
MAXIM-2S (Tu et al., 2022) MLP 38.06 0.977 22.50 80.13
SFNet (Cui et al., 2022) CNN 38.21 0.974 13.27 70.00
DRCNet (Li et al., 2022b) CNN 38.23 0.976 18.90 -
Restormer (Zamir et al., 2022) Transformer 38.99 0.978 26.13 92.85
LIR (Ours) CNN 38.96 0.983 7.31 62.51
3.3.1. ResBlock and ResCABlock
@@
Both ResBlock and ResCABlock are up-down style residual blocks that include a learnable parameter in the trunk. The only difference between ResBlock and ResCABlock is that the latter has a channel attention module at the tail. The Channel Attention module (Zhang et al., 2018a), based on Squeeze-and-Excitation (SE) (Hu et al., 2018), has gained much popularity for its ability to extract channel-wise information. Many studies (Li et al., 2023; Chen et al., 2022b, a; Zhao et al., 2020) have directly utilized this module in their models. Following them, in our Attention Block, we also incorporate the Channel Attention module into the ResBlock to extract channel-wise information. In addition, the structure of ResBlock and ResCABlock are designed to be simple to meet lightweight and efficient requirements.
Figure 6. The visual comparisons for LIR, Restormer (Zamir et al., 2022), AirNet (Li et al., 2022a), and MPRNet (Zamir et al., 2021). LIR produces state-of-the-art visual results.
3.3.2. Patch Attention
@@
To give the Attention Block the ability to capture global information, we introduce the Patch Attention (PA) (Figure 4). In the self-attention part of the transformer, the attention map is calculated using two matrix multiplications and a softmax operation between q, k, and v, which incurs significant computational overhead. Instead, we simplify the attention map computation in our PA by using just one MLP and one element-wise multiplication.
@@
Figure 7. The visual comparisons for LIR, FFA-Net (Qin et al., 2020), GridDehazeNet (Liu et al., 2019), and AirNet (Li et al., 2022a). LIR produces better visual results.
@@
Table 2. Image denoising results evaluated by PSNR on CBSD68 and Urban100 datasets. Where the background of light blue for the method that is learned to handle various noise levels.
Method Type CBSD68 (Martin et al., 2001) Urban100 (Huang et al., 2015) Params [M] GFLOPs
$\sigma=15$ $\sigma=25$ $\sigma=50$ $\sigma=15$ $\sigma=25$ $\sigma=50$
FD-GAN (Dong et al., 2020a) GAN 30.25 28.81 26.43 - - - 13.98 38.90
DL (Fan et al., 2019) CNN 33.05 30.41 26.90 - - - - -
MPRNet (Zamir et al., 2021) CNN 33.54 30.89 27.56 - - - 20.13 960.39
FFDNet (Zhang et al., 2018b) CNN 33.87 31.21 27.96 33.83 31.40 28.05 - -
DnCNN (Zhang et al., 2017) CNN 33.90 31.24 27.95 32.98 30.81 27.59 - -
AirNet (Li et al., 2022a) CNN 33.92 31.26 28.00 34.40 32.10 28.88 8.93 169.47
BRDNet (Tian et al., 2020) CNN 34.10 31.43 28.16 34.42 31.99 28.56 - -
IDR (Zhang et al., 2023b) CNN 34.11 31.60 28.14 33.82 31.29 28.07 15.34 -
DRUNet (Zhang et al., 2021a) CNN 34.30 31.69 28.51 34.81 32.60 29.61 32.64 80.78
Restormer (Chen et al., 2022a) Transformer 34.39 31.78 28.59 35.06 32.91 30.02 26.13 92.85
LIR (Ours) CNN 34.31 31.69 28.50 34.81 32.59 29.57 7.31 62.51
4. Experiment
4.1. Experiment Settings
We set the number of LAA Blocks in the LIR to [3,3,3,3,4], and the width to 48. During the training, we use the $L_{1}$ loss function, AdamW ($\beta_{1}$=0.9, $\beta_{2}$=0.999) optimizer with an initial learning rate of $3e^{-4}$, and set weight decay as $1e^{-4}$. Random image cropping and horizontal flipping are used as data augmentation strategies. Following (Zamir et al., 2022), we use the cosine annealing strategy to gradually decrease the learning rate from $3e^{-4}$ to $1e^{-6}$, and use the progressive learning, where training patch size and batch size pairs are: [($128^{2}$,32), ($160^{2}$,16), ($256^{2}$,8)] at iterations [250K, 250K, 250K] for deblurring, deraining, and denoising. For dehazing, we set [($128^{2}$,32), ($160^{2}$,16)] at iterations [350K, 400K]. The number of parameters and computation for each method are calculated based on the input size of (3,192,192). Specifically, parameters are counted using the PyTorch, and the computation is measured using the thop package.
4.2. Image Deraining
The Rain14000 (Fu et al., 2017) (11,200 training images) and Rain100H (Yang et al., 2017) (1,254 training images) are used to train LIR, and the Rain100L (Yang et al., 2017) datasets is used for evaluation. The results are presented in Table 1, where LIR outperforms the DRCNet (Li et al., 2022b) and achieves comparable performance to Restormer (Zamir et al., 2022) with only a slight 0.02 dB decrease on PSNR with minimal parameters. In addition, our LIR achieves state-of-the-art on SSIM, which significantly outperforms Restormer. It is worth noting that with fewer parameters and computations, LIR significantly outperforms SFNet (Cui et al., 2022), the same lightweight design method as ours. Figure 5 shows visual comparisons among LIR, Restormer, SFNet, and HINet. It can be demonstrated that LIR produces state-of-the-art visual results, despite LIR having a slightly lower PSNR than Restormer.
Table 3. Image dehazing results evaluated by PSNR and SSIM on SOTS outdoor dataset.
Method Type SOTS outdoor (Li et al., 2018) Params [M] GFLOPs
PSNR SSIM
FD-GAN (Dong et al., 2020a) GAN 23.15 0.920 13.98 38.90
AirNet (Li et al., 2022a) CNN 23.18 0.900 8.93 169.47
IDR (Zhang et al., 2023b) CNN 25.24 0.943 15.34 -
GCANet (Chen et al., 2019) CNN 30.23 0.980 - -
GridDehazeNet (Liu et al., 2019) CNN 30.86 0.982 0.96 13.69
U2-Former (Ji et al., 2021) Transformer 31.10 0.976 - -
FFA-Net (Qin et al., 2020) CNN 33.38 0.980 4.46 161.74
MSBDN (Dong et al., 2020b) CNN 33.67 0.985 - -
MAXIM-2S (Tu et al., 2022) MLP 34.19 0.985 22.5 80.13
LIR (Ours) CNN 34.85 0.985 7.31 62.51
@@ 4.2. Image Deraining
Table 4. Image deblurring results evaluated by PSNR and SSIM on GoPro and HIDE datasets.
Method Type GoPro (Nah et al., 2017) HIDE (Shen et al., 2019) Params [M] GFLOPs
PSNR SSIM PSNR SSIM
IDR (Zhang et al., 2023b) CNN 27.87 0.846 - - 15.34 -
DeblurGAN-v2 (Kupyn et al., 2019) GAN 29.55 0.934 26.61 0.875 59.16 20.12
SRN (Tao et al., 2018) RNN 30.26 0.934 28.36 0.915 - -
Gao et al. (Gao et al., 2019) CNN 30.90 0.935 29.11 0.913 - -
DBGAN (Zhang et al., 2020) GAN 31.10 0.942 28.94 0.915 11.58 189.71
MT-RNN (Park et al., 2020) RNN 31.15 0.945 29.15 0.918 - -
DMPHN (Zhang et al., 2019) CNN 31.20 0.940 29.09 0.924 21.70 -
Zhe et al. (Jiang et al., 2020) CNN 31.79 0.949 - - - -
Suin et al. (Suin et al., 2020) CNN 31.85 0.948 29.98 0.939 - -
SPAIR (Purohit et al., 2021) CNN 32.06 0.950 30.29 0.931 - -
LIR (Ours) CNN 32.19 0.931 30.31 0.917 7.31 62.51
4.3. Image Dehazing
The OTS (Li et al., 2018) (72,315 training images) is used to train LIR, and the Synthetic Objective Testing Set (SOTS) outdoor (Li et al., 2018) is used for evaluation. In the SOTS outdoor results (Table 3), LIR significantly outperforms both the MAXIM-2S (Tu et al., 2022) (MLP-based) and the U2-Former (Ji et al., 2021) (transformer-based) with improvement in PSNR of 0.66 dB and 3.75 dB. In addition, LIR achieves comparable performance on PSNR and better performance on SSIM to DehazeFormer. Visual comparisons are shown in the Figure 7.
4.4. Image Denoising
We perform denoising experiments on synthetic benchmark datasets generated by additive white Gaussian noise. The DFWB (DIV2K (Agustsson and Timofte, 2017), Flickr2K, WED (Ma et al., 2016), BSD500 (Arbelaez et al., 2010)) datasets with 8,594 training images are used to train LIR. For evaluation, we use the CBSD68 (Martin et al., 2001) and Urban100 (Huang et al., 2015) datasets. The background of the light blue color in the Table 2 indicates that the method is trained using a blind denoising setting (random $\sigma$ from [0,50]) and then evaluated under a fixed $\sigma$ (0, 25, or 50). Other methods are trained with a fixed $\sigma$ and evaluated with the same $\sigma$ as training. As shown in the Table 2, LIR achieves comparable performance to Restormer (Zamir et al., 2022) and significantly outperforms all other models on the CBSD68 dataset with a blind denoising setting. Moreover, it is worth noting that with a 77% reduction in parameters and a 20% reduction in computations, our LIR achieves comparable performance to DRUNet (Zhang et al., 2021a) on Urban100, while the DRUNet is trained under fixed $\sigma$ setting and our LIR is under random $\sigma$ setting. The visual comparisons between LIR, Restormer, AirNet, and MPRNet are shown in Figure 6. Although LIR is not as high as Restormer on PSNR, the better visual result is achieved.
4.5. Image Deblurring
@@
The GoPro (Nah et al., 2017) (2,103 training images) dataset is used to train LIR, and the GoPro (1,111 images) and HIDE (Shen et al., 2019) (2,025 images) datasets are used for evaluation. Table 4 demonstrates that our LIR significantly outperforms other methods with fewer parameters and computations. Specifically, LIR is 0.13 dB higher than SPAIR (Purohit et al., 2021), and 0.34 dB significantly higher than Suin et al (Suin et al., 2020).
4.6. Ablation Studies
@@
We conduct ablation experiments on the deraining task using different types. Type 1 refers to LIR, Type 2 refers to using vanilla global and local residual connections in the model without removing the degradation in residual connections through structural design, Type 3 refers to no Patch Attention module, Type 4 refers to no Adaptive Filter, and Type 5 refers to using vanilla global and local residual connections in the model and not include Patch Attention module and Adaptive Filter. The results in the Table 5 demonstrate that the absence of the Adaptive Filter has the greatest negative impact on LIR, where PSNR is droped by 0.3 dB. Furthermore, it also shown that our proposed Patch Attention module is efficient and computation-friendly, only bringing 0.8 M parameters and 1.81 GFLOPs computations. Although Type 1 only improved performance by 0.09 dB compared to Type 2, structural design for residual connections is the key to the final visual improvement.
In addition, we present the intermediate feature maps of Type 1 and Type 4 (Figure 9). Obviously, LIR with Adaptive Filter has a better visual result than without Adaptive Filter, with less noise and rain trails. We also present the intermediate feature maps of Type 1 and Type 2 (Figure 8), where the former has a better visual result than the latter, with fewer rain trails and better contour. More visual results about Type 1, Type 2, and Type 5 please see Supplementary Materials.
Table 5. Ablation studies are performed on Rain100L dataset. Where Vanilla Con. (Connections) refers to using vanilla global and local residual connections in the model without removing the degradation in residual connections.
Type 1 Type 2 Type 3 Type 4 Type 5
Vanilla Con.
Adaptive Filter
Patch Attention
Params [M] 7.31 6.86 7.31 6.51 5.88
GFLOPs 62.51 49.15 62.51 62.15 49.01
PSNR 38.96 38.87 38.66 38.85 37.65
5. Conclusion and Limitation
Figure 8. The visualization of intermediate feature maps of the Type 1 (top) and the Type 2 (bottom) in the deraining task.
Figure 9. The visualization of intermediate feature maps of the Type 1 (top) and the Type 4 (bottom) in the deraining task.
5.1. Conclusion